# Beyond Needle(s) in the Embodied Haystack: Environment, Architecture, and Training Considerations for Long Context Reasoning

### v2

@@ B.1 Interactive Evaluation in $\infty$-Thor
16:return Total accumulated reward $R$
Appendix C Additional Results
Figure 7: Agent’s reward across different experimental configurations: (a) context extension methods at fixed scaling (x4), (b) varying YaRN scaling factors, (c) fine-tuning with different context lengths using Context Parallelism, (d-e) combinations of scaling during both training and inference, and (f) summary of the most effective strategies.
C.1 Interactive Evaluation: High-level Planning
Results and Discussion. Figure 7 presents the accumulated rewards over time across six experimental configurations. The Planner trajectory represents the performance upper bound. Our analysis focuses on addressing the following key questions:
@@ C.1 Interactive Evaluation: High-level Planning
Q. Does combining context extension methods during both training and inference provide additional benefits? Results of experiments with scaling at both training and evaluation (Figures 7(d) and 7(e) in Appendix) indicate that additional scaling at evaluation after fine-tuning with scaled RoPE provides no further performance improvement and may degrade performance at shorter context lengths ($\leq$300K tokens).
Based on these observations, we can conclude that fine-tuning strategies are most effective when long-trajectory datasets are available. In the absence of extensive training data, employing YaRN scaling at x4 yields performance comparable to the Planner upper-bound, particularly within context lengths under 200K tokens (Figure 7(f)).
C.2 Interactive Evaluation: Low-level Manipulation
Table 6: Success rates of OpenVLA-7B and SpatialVLA-4B on low-level Pick-up and Put tasks in ManipulaTHOR.
Task OpenVLA-7B SpatialVLA-4B
Pick-up 16.52% 18.70%
Put 2.08% 1.19%
We evaluate the ability of existing VLA models, OpenVLA-7B (Kim et al., 2024) and SpatialVLA-4B (Qu et al., 2025), to control robot arms on low-level manipulation tasks. The evaluation protocol follows the procedure described in Appendix B.1, with the only difference being that low-level VLA models are used to directly execute Pick-up and Put actions through arm control. For the text input, we provide task-specific instructions (“Pick-up” or “Put” object), since the existing models are only trained on single-task formulations. For the image input, we use an ego-centric camera view. Since these VLA models were originally trained on datasets with different viewing angles, there remains considerable room for improvement by incorporating additional camera perspectives during training and evaluation.
@@ C.2 Interactive Evaluation: Low-level Manipulation
Figure 6 presents accumulated rewards on low-level manipulation tasks. Both models underperform relative to the Planner baseline, largely due to differences in robot arm configuration and the out-of-distribution nature of the visual inputs. The results show that SpatialVLA achieves slightly higher and more stable rewards than OpenVLA across long sequences.
Table 6 reports task success rates. Overall performance remains low, with both models struggling on manipulation tasks. We attribute these results primarily to two factors: (1) ManipulaTHOR’s egocentric view, which provides nearly identical images for fine-grained arm movements $(\Delta x,\Delta y,\Delta z)$ smaller than 0.01, making it difficult for VLA models to perceive subtle adjustments; and (2) discrepancies between the robot arm in ManipulaTHOR and those assumed by the VLA models, causing many predicted actions to fail in the physical simulation. We expect that fine-tuning on our dataset could provide an effective solution to bridge these gaps.
Appendix D Limitations
While $\infty$-Thor enables the generation of arbitrarily long trajectories, the diversity of environment layouts in AI2-THOR is inherently limited. This can lead to repetitive agent behaviors within certain scenes. For instance, compared to kitchen and living room scenes, bedroom scenes tend to involve fewer action types, mostly constrained to simple pick-and-place tasks within small spatial areas. Due to the limited scene size and low task diversity, we excluded bathroom scenes from our dev and test sets. In future work, we plan to integrate $\infty$-Thor with ProcTHOR (Deitke et al., 2022), which supports procedurally generated environments, enabling a broader range of dynamic and diverse scene configurations.
Additionally, although the GT action sequences generated by the PDDL-based planner are sufficient for task completion, they are not guaranteed to be optimal. This may lead to agents learning suboptimal behaviors when trained solely on these demonstrations. Incorporating learning from exploration or reinforcement-based optimization could improve policy quality.
Finally, inference with LLMs in long-context settings remains a computational bottleneck, particularly as context lengths approach 1M tokens. Since our long-horizon tasks require access to information spread throughout the entire trajectory, full-context inference becomes increasingly expensive, even with Context Parallelism. One promising direction is to equip agents with memory systems that selectively keep relevant information from prior steps, allowing the model to reason without reprocessing the full context at each step. Other architectural solutions, such as sparse attention mechanisms (Han et al., 2024; Jiang et al., 2024) or state-space models with linear-time inference (Gu & Dao, 2024), also hold potential for scalable long-context reasoning.

### v3

@@ C.2 Online Evaluation: High-level Planning
16:
17:  return Total accumulated reward $R$
Results and Discussion. Figure 8 presents the accumulated rewards over time across four experimental configurations. The Planner trajectory represents the performance upper bound. Our analysis focuses on addressing the following key questions:
@@
Q. Does combining context extension methods during both training and inference provide additional benefits? Results of experiments with scaling at both training and evaluation (Figures 8(d)) indicate that additional scaling at evaluation after fine-tuning with scaled RoPE provides no further performance improvement and may degrade performance at shorter context lengths ($\leq$300K tokens).
Based on these observations, we can conclude that fine-tuning strategies are most effective when long-trajectory datasets are available. In the absence of extensive training data, employing YaRN scaling at x4 yields performance comparable to the Planner upper-bound, particularly within context lengths under 200K tokens.
C.3 Online Evaluation: Low-level Manipulation
We evaluate the ability of existing VLA models, OpenVLA-7B (Kim et al., 2024) and SpatialVLA-4B (Qu et al., 2025), to control robot arms on low-level manipulation tasks. The evaluation protocol follows the Algorithm 2, with the only difference being that low-level VLA models are used to directly execute Pick up and Put actions through arm control. For the text input, we provide task-specific instructions (Pick up or Put object), since the existing models are only trained on single-task formulations. For the image input, we use an ego-centric camera view. Since these VLA models were originally trained on datasets with different viewing angles, there remains considerable room for improvement by incorporating additional camera perspectives during training and evaluation.
@@ C.3 Online Evaluation: Low-level Manipulation
Result. Figure 9 presents accumulated rewards on low-level manipulation tasks. Both models underperform relative to the Planner baseline, largely due to differences in robot arm configuration and the out-of-distribution nature of the visual inputs. The results show that SpatialVLA achieves slightly higher and more stable rewards than OpenVLA across long sequences.